Polynomial P(x) contains only terms of odd degree. when P(x) is divided by (x-3),the remainder is 6.If P(x) is divided by(x2-9) then remainder is g(x).Find the value of g(2).

Grade:12

1 Answers

harshit goswami
13 years ago

i hav assumed the odd degree polynomial p(x)= ax3 + bx, then on dividing it by (x - 3) we get remainder= 3(9a + b) also on dividing it by x2 - 9 we get g(x)=x(9a + b) on solving both we get g(2)=4.